# Toward Low-Latency, High-Fidelity Calibration of the LIGO Detectors with Enhanced Monitoring Tools

### v1

@@ Appendix C Limits on Estimation of Systematic Error on Calibrated Strain
Section 2.4 introduced historical and current methods used to estimate the systematic error on the detector response $\eta_{R}$ (see (7)). The primary limit on the instantaneous estimate of $\eta_{R}^{\textrm{(meas)}}$ has historically come from the uncertainty of the GPR fit. [^4] The superscript (meas) refers to the measurement of a given quantity. We can measure the error, $\eta_{R;A_{i}}^{\textrm{(meas)}}$ and $\eta_{R;C}^{\textrm{(meas)}}$, at any given time with the accuracy and precision limited by the absolute reference system and our knowledge of the well-known parameters. It is the model of those errors that has high uncertainty, for several reasons. First, the model is informed by many repeated measurements that are taken outside of the time in question. Second, those measurements are conditioned before fitting with well-known components (e.g. electronics) divided out. Third, the GPR is made with an idealistic kernel of hyperparameters that assumes little-to-no frequency-dependent systematic error, rather than the reality of a multitude of frequency-dependent features. Fourth, we tune the bounds of the hyperparameters to cover all systematic error possibilities within the epoch, resulting in an uncertainty that is conservatively high. Finally, the GPR fit is propagated through the loop topology of the response function, which in the most sensitive region of the detector amplifies any errors that may have been in the original GPR fits. For these reasons, the claimed uncertainty in $\eta_{R}$ for any instant in time, let alone for the entire epoch, is likely several factors larger than the underlying true systematic error at any given time. Figure 15 in [12] shows this well; the uncertainty at each measured frequency point is barely visible on the scale, where the modeled systematic error is an appreciable portion of the plot.

### v2

@@ Appendix C Limits on Estimation of Systematic Error on Calibrated Strain
Section 2.4 introduced historical and current methods used to estimate the systematic error on the detector response $\eta_{R}$ (see (7)). The primary limit on the instantaneous estimate of $\eta_{R}^{\textrm{(meas)}}$ has historically come from the uncertainty of the GPR fit [^4] The superscript (meas) refers to the measurement of a given quantity.. We can measure the error, $\eta_{R;A_{i}}^{\textrm{(meas)}}$ and $\eta_{R;C}^{\textrm{(meas)}}$, at any given time with the accuracy and precision limited by the absolute reference system and our knowledge of the well-known parameters. It is the model of those errors that has high uncertainty, for several reasons. First, the model is informed by many repeated measurements that are taken outside of the time in question. Second, those measurements are conditioned before fitting with well-known components (e.g. electronics) divided out. Third, the GPR is made with an idealistic kernel of hyperparameters that assumes little-to-no frequency-dependent systematic error, rather than the reality of a multitude of frequency-dependent features. Fourth, we tune the bounds of the hyperparameters to cover all systematic error possibilities within the epoch, resulting in an uncertainty that is conservatively high. Finally, the GPR fit is propagated through the loop topology of the response function, which in the most sensitive region of the detector amplifies any errors that may have been in the original GPR fits. For these reasons, the claimed uncertainty in $\eta_{R}$ for any instant in time, let alone for the entire epoch, is likely several factors larger than the underlying true systematic error at any given time. Figure 15 in [12] shows this well; the uncertainty at each measured frequency point is barely visible on the scale, where the modeled systematic error is an appreciable portion of the plot.